OpenStudy (anonymous):

if twice a number added to 3 is divided by the number plus 2, the result is five thirds. find the number??

OpenStudy (anonymous):

"If twice a numer" (2x) "added to 3" (2x+3) "is divided by the number plus 2" (2x+3)/(x+2), "the result is five thirds" So, in the ends you have \[\frac{ 2x+3 }{ x+2 }=\frac{ 5 }{ 3 }\], and you solve for x, got it?

OpenStudy (anonymous):

x=1?

OpenStudy (anonymous):

Yes, that's correct